The larger of two numbers is 1 less the 3 times the smaller. The sum is 63. Find the numbers
Answer by doukungfoo(195)

Lets call the smaller number x
Given: the larger number is 1 less than 3 times the smaller number.
So if the smaller number = x
then
the larger number = 3x-1
Given: their sum is 63
so
Smaller number plus larger number equals 63
write an equation and solve for x
x + 3x - 1 = 63
4x - 1 = 63
4x = 63 + 1
4x = 64
x = 16
Answer: smaller number is 16
use this to find the larger number
larger number = 3x-1
larger number = 3(16)-1
larger number = 48 - 1
larger number = 47
Answer: larger number is 47
Check answers:
their sum should equal 63
16 + 47 = 63
63 = 63